# Efficient Data-driven Scene Simulation using Robotic Surgery Videos via Physics-embedded 3D Gaussians

### v2

@@ 1 Introduction
Surgical scene 3D reconstruction has been studied in recent works such as EndoNeRF [25], EndoSurf[31], LerPlane[29] and ForPlane[28]. These NeRF-based methods can be used to derive simulatable representation, but implementing this typically requires sophisticated post-processing [24, 26] for tetrahedral mesh convert and shape representation transition [15, 18]. Therefore, using NeRF-based reconstruction results for simulation is suboptimal in practice. Very recently, Gaussian Splatting (GS) [9] has emerged as a promising alternative to NeRF[17], offering superior 3D reconstruction results and faster inference speed overall. Several concurrent works [32, 12, 8] have applied 3D-GS for surgical reconstruction. Unlike NeRF, 3D-GS uses explicit Gaussian representation, making it more suitable for shape editing[2], deformation tracking[16] and simulation[27]. Notably, a successful simulation from 3D-GS relies on clean Gaussians generated from the data, which requires the GS-based reconstruction framework to be particularly designed. For instance, EndoGS[32] is not suitable for simulation purpose, because it tends to generate floating or slim Gaussians in order to fit images, which would lead to noticeable artifacts in simulations (see Fig. 3 in Sec. 3).
@@
In this paper, we endeavor to simulate reconstructed surgical scenes captured from single-viewpoint stereo endoscope in a completely data-driven manner (called SimEndoGS). We embark on integrating the emerging 3D Gaussian Splatting and MPM framework to tackle surgical scene simulation. We summarize our contributions as follows: 1) An efficient and automated pipeline consisting of designed 3D-GS reconstruction module and a subsequent efficient MPM simulation module. 2) A geometric regularization is proposed to overcome overfitting issue of Gaussian Splatting on endoscopic data. 3) The Neo-Hookean model and an adapted MPM solver for 3D Gaussian are seamlessly integrated into our pipeline to perform physics-embedded soft tissue simulation. Experimental results on our robotic surgery videos demonstrate the success of our data-driven reconstruction and simulation pipeline, which can support soft-tissue interactions with high fidelity and approaching real-time speed.
@@ 2.3 Geometrically Regularized Optimization for 3D Gaussian
In endoscopic scene, the limited movement range of camera makes it difficult for Gaussian Splatting to learn the geometry information and then lead to overfitting. To overcome this problem, we initialize the positions of Gaussians by reprojecting the depth maps $\{\bm{D}_{i}\}^{T}_{i=1}$ to get a reasonable geometric structure before the training. During the optimization, we add the Huber loss between estimated depth maps and rendered depth maps to the objective function of original Gaussian Splatting framework. The depth value of 3D Gaussian at pixel $\mathbf{p}$ shares the same computation pattern with Eq. 1. Our complete loss function is:
@@ 2.4 Physically-based 3D Gaussians Simulation
where $\mathbf{F}$ is the deformation gradient obtained from MPM solver, $\mathbf{\Sigma}$ is the initial covariance matrix and $\mathbf{{\Sigma}^{\prime}}$ is the updated covariance matrix. We utilize Neo-Hookean[21] constitutive model in MPM to predict tissue deformation because of its simplicity and computation efficiency. The first Piola-Kirchoff stress of Neo-Hookean model, denoted as $\mathbf{PK_{1}}$, is computed as follows:
@@
By adjusting the values of Young’s modulus $E$ and Poisson ratio $\nu$, we can control the physical behavior of the endoscopic tissue, such as its stiffness, in a physically interpretable way.
@@ 3 Experiments
We extensively evaluated our proposed data-driven surgical simulation on five different endoscopic scenes, using corresponding stereo videos collected from the publicly available datasets [25, 30] or our in-house data. These videos encompassed common surgical procedures involving the manipulation, pushing and pulling of deformable tissues. We compared our method with an GS-based baseline. Similar to [32, 12, 8], we quantitatively evaluated the surgical scene reconstruction using PSNR. For the simulation performance, since it is difficult for a quantitative evaluation, we reported the simulation efficiency and qualitatively measured the simulation performance by comparing the texture and geometry details under different simulation interactions. We also conducted an ablation study to evaluate the proposed depth supervision module, anisotropy regularization module and Gaussian padding module through qualitative comparison.

### v3

# SimEndoGS: Efficient Data-driven Scene Simulation using Robotic Surgery Videos via Physics-embedded 3D Gaussians

@@ 1 Introduction
Dynamic surgical scene reconstruction has been studied in recent works such as EndoNeRF [23], EndoSurf[29], LerPlane[27] and ForPlane[26]. All these NeRF-based methods typically requires sophisticated post-processing [22, 24] to turn implicit representation to simulatable mesh [14, 16]. This transformation increases the time cost and lose the reconstruction detail of surgical scenes, making NeRF-based methods suboptimal for simulation in practice. Very recently, Gaussian Splatting (GS) [8] has emerged as a promising alternative to NeRF[15], offering superior 3D reconstruction results and faster inference speed overall. Several concurrent works [30, 11, 7] have applied 3D-GS for surgical reconstruction. Unlike NeRF, 3D-GS uses explicit Gaussian representation, making it feasible for physical simulation without any modality transformation[25]. It should be emphasized that all dynamic surgical scene reconstruction methods mentioned above model the tissue deformation using neural networks which take time and spatial location as input. Due to the lack of consideration for material properties and physical rules, all these learning-based methods could not support interactive surgical simulation.
@@
In this paper, we endeavor to simulate reconstructed surgical scenes captured from single-viewpoint stereo endoscope in a completely data-driven manner (called SimEndoGS). To the best of our knowledge, this is the first work to utilize the emerging 3D Gaussian Splatting and Material Point Method (MPM) framework to address the challenge of surgical scene simulation. We summarize our contributions as follows: 1) An efficient and automated pipeline consisting of designed 3D-GS reconstruction module and a subsequent efficient MPM simulation module. 2) A geometric regularization is proposed to overcome overfitting issue of Gaussian Splatting on endoscopic data. 3) The Neo-Hookean model and an adapted MPM solver for 3D Gaussian are seamlessly integrated into our pipeline to generate physically-based deformation in surgical scenes. Experimental results on our robotic surgery videos demonstrate the success of our data-driven reconstruction and simulation pipeline, which can support soft-tissue interactions with realistic deformation and approaching real-time speed.
@@ 2.3 Geometrically Regularized Optimization for 3D Gaussian
In endoscopic scene, the limited movement range of camera makes it difficult for GS to learn the geometry information and then lead to overfitting. To overcome this problem, we initialize the positions of Gaussians by reprojecting the depth maps $\{\bm{D}_{i}\}^{T}_{i=1}$ to get a reasonable geometric structure before the training. During the optimization, we add the Huber loss between estimated depth maps and rendered depth maps to the objective function of original Gaussian Splatting framework. The depth value of 3D Gaussian at pixel $\mathbf{p}$ shares the same computation pattern with Eq. 1. Our complete loss function is:
@@ 2.4 Physically-based 3D Gaussians Simulation
where $\mathbf{F}$ is the deformation gradient obtained from MPM solver, $\mathbf{\Sigma}$ is the initial covariance matrix and $\mathbf{{\Sigma}^{\prime}}$ is the updated covariance matrix. We utilize Neo-Hookean[19] constitutive model in MPM to model the nonlinear stress-strain relationship of biological tissue. The first Piola-Kirchoff stress of Neo-Hookean model, denoted as $\mathbf{PK_{1}}$, is computed as follows:
@@
$\mathbf{PK_{1}}$ will participate in the following computation of grid forces and then influence the tissue dynamic. By adjusting the values of Young’s modulus $E$ and Poisson ratio $\nu$, we can control the physical behavior of the endoscopic tissue, such as its stiffness, in a physically interpretable way.
@@ 3 Experiments
We extensively evaluated our proposed data-driven surgical simulation on five different endoscopic scenes, using stereo endoscopic videos introduced in[28], EndoNeRF datasets [23] and our in-house data. These videos encompassed common surgical procedures involving the manipulation, pushing and pulling of deformable tissues. We compared our method with an GS-based baseline. Similar to [30, 11, 7], we quantitatively evaluated the surgical scene reconstruction using PSNR. For the simulation performance, since it is difficult for a quantitative evaluation, we reported the simulation efficiency and qualitatively measured the simulation performance by comparing the texture and geometry details under different simulation interactions. We also conducted an ablation study to evaluate the proposed depth supervision module, anisotropy regularization module and Gaussian padding module through qualitative comparison.